Amy measures the slant height of a cone-shaped cup and finds that it is 12 cm. The height is 10 cm. Determine the volume of water in the cup if Amy fills it to the top.


1
Expert's answer
2020-11-09T15:40:50-0500

Explanations & Calculations


  • If the radius of the base of the cone (the circle) is r, then considering a cross section of the cone through a diameter of the base circle, you can sense of a usage of Pythagoras theorem to find the radius (r) of the base.
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small r^2+(10cm)^2 &= \small (12cm)^2\\
\small r&= \small \sqrt{12^2-10^2}\\
&= \small \bold{2\sqrt{11} cm}
\end{aligned}"

  • And the volume of a cone is given is

"\qquad\qquad
\begin{aligned}
\small V &= \small \frac{1}{3}\pi r^2h\\
\small 
\end{aligned}"

  • Therefore knowing both r & h, the volume is

"\qquad\qquad
\begin{aligned}
\small V&= \small \frac{1}{3}\times\pi \times(2\sqrt{11}cm)^2\times10cm\\
\small &= \small \bold{460.767cm^3}
\end{aligned}"